# Sparse Autoencoders for Sequential Recommendation Models: Interpretation and Flexible Control

### v1

@@ Abstract.
Many current state-of-the-art models for sequential recommendations are based on transformer architectures. Interpretation and explanation of such black box models is an important research question, as a better understanding of their internals can help understand, influence, and control their behavior, which is very important in a variety of real-world applications. Recently sparse autoencoders (SAE) have been shown to be a promising unsupervised approach for extracting interpretable features from language models. These autoencoders learn to reconstruct hidden states of the transformer’s internal layers from sparse linear combinations of directions in their activation space.
This paper is focused on the application of SAE to the sequential recommendation domain. We show that this approach can be successfully applied to the transformer trained on a sequential recommendation task: learned directions turn out to be more interpretable and monosemantic than the original hidden state dimensions. Moreover, we demonstrate that the features learned by SAE can be used to effectively and flexibly control the model’s behavior, providing end-users with a straightforward method to adjust their recommendations to different custom scenarios and contexts.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 19th ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Sequential recommender systems capture the dynamic nature of user preferences by taking into account the temporal order of user actions. This approach is particularly valuable in domains where user interests evolve and change rapidly, so the sequence of events matters a lot. By incorporating this sequential information, models can provide more accurate and relevant recommendations.
@@
Transformer-based models, such as SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and their numerous improvements, can effectively model complex sequential patterns. Interpretation and explanation of these models are of high practical importance due to their increasing usage both in research and industrial applications. By gaining insights into their internal workings, researchers and developers can better debug the models and identify potential biases and issues, which can lead to improvements in model design and performance. Explainable recommendations, providing transparency into why certain recommendations are made, can improve user trust and the reliability of recommendations. Finally, interpretability may help to influence and adjust recommendations to specific contexts, individual user preferences, and needs, leading to a more satisfying and personalized user experience.
The idea that neural networks can organize information in interpretable patterns has guided research for decades, supported by evidence for the linear representation of interpretable concepts (Smolensky, 1986; Marks and Tegmark, 2023; Nanda et al., 2023; Elhage et al., 2022; Park et al., 2024), formalized as the Linear Representation Hypothesis (LRH). While the complete picture goes beyond LRH (Csordás et al., 2024; Engels et al., 2024a, b; Smith, 2024), its “weak” form has nonetheless demonstrated significant utility. However, identifying “meaningful” linear subspaces presents challenges, as focusing on individual neurons has proven ineffective; single neurons typically encode multiple concepts – a phenomenon referred to as polysemanticity (Elhage et al., 2022; Bolukbasi et al., 2021). To address this, training Sparse AutoEncoders (SAE) on the hidden states of models has emerged as one of the most effective methods for extracting more monosemantic features (i.e., neurons of the SAE’s decoder). So far, this approach has been mostly applied to Large Language Models (LLM) and Vision-Language Models (VLM)(Cunningham et al., 2023; Makelov et al., 2024; Gao et al., 2024; Templeton et al., 2024), with some examples in the vision (Gorton, 2024; Le et al., 2024; Daujotas, 2024) and other domains (He et al., 2024).
In this work, we apply SAE to sequential recommendations. Evaluating the features learned by an SAE remains challenging, particularly in the text domain, where the absence of explicit, well-defined attributes makes it difficult to measure their relevance (Makelov et al., 2024). The field of recommendation systems, on the other hand, is particularly well-suited for applying this approach because the items are usually associated with explicit attributes, providing a clearer basis for assessing the usefulness of the learned features.
Previous works also show that steering – the process of adding vectors that represent SAE features to a model’s hidden state – can often be used to control the model’s outputs in a way that aligns with the interpreted meaning of the SAE feature  (Templeton et al., 2024). We show that this property holds for recommendation models, allowing the desired properties to be more or less present in the recommendations.
@@
  • We extend the sparse autoencoders for sequential recommendations: propose the ways to find interpretable features and metrics to assess the interpretability of SAE representations.
  • We demonstrate that SAE learns a lot of interpretable, non-trivial and semantically distinct features.
  • We show that SAE allows to achieve flexible control over model predictions for better alignment with users specific needs and contexts.
@@ 2. Related Work
Various approaches have been proposed to interpret recommendation models. In the study (Tsang et al., 2020), automatic detection of important feature interactions in neural networks is introduced. Other interpretation techniques are based on language modeling (Hada et al., 2021; Li et al., 2021, 2023b), attention mechanism (Xiao et al., 2017; Tal et al., 2021), counterfactual methods (Tan et al., 2021; Kaffes et al., 2021), and other explanation strategies. While these methods help reveal aspects of the model’s internal structure, they are often closely tied to specific architectures, which limits their generalizability.
@@ 3. Sparse Autoencoders
SAE has emerged as a tool in the field of LLM mechanistic interpretability through the employment of techniques from sparse dictionary learning that were previously applied to tasks such as understanding neuronal signals in the visual cortex  (Olshausen and Field, 1997, 2004; Cunningham et al., 2023). Building on experiments with smaller transformer models (Elhage et al., 2022; Cunningham et al., 2023), SAE-based approaches have been further developed and scaled to larger models, including cutting-edge proprietary LLMs. These advancements continue to provide insights into model internals and offer promising strategies for better controlling model generation (Templeton et al., 2024; Gao et al., 2024; Lindsey et al., 2024).
Sparse autoencoders (SAEs) are a class of autoencoders designed to produce compact and interpretable latent representations by ensuring that only a small subset of hidden units becomes active for any given input. In the classic formulation, an input $x\in\mathbb{R}^{n}$ is mapped to a hidden representation $h(x)\in\mathbb{R}^{m}$ through an encoder and then reconstructed by a decoder to approximate the original input, $\hat{x}$. This can be expressed as $h(x)=\sigma(Wx+b)$ and $\hat{x}=W^{\prime}h(x)+b^{\prime}$, where $W,W^{\prime}$ are the weight matrices (often tied, i.e., $W^{\prime}=W^{\top}$), $b,b^{\prime}$ are bias terms, and $\sigma(\cdot)$ is a nonlinear activation function such as ReLU. The training objective is to minimize the reconstruction loss, $\|x-\hat{x}\|_{2}^{2}$ while imposing a sparsity constraint on the hidden representation $h(x)$.
A common approach to enforce sparsity on SAE is to add an $L_{1}$ regularization term on $h(x)$, resulting in the following objective function  (Sharkey et al., 2022; Rajamanoharan et al., 2024a): $L(x)=\|x-\hat{x}\|_{2}^{2}+\lambda\|h(x)\|_{1}$, where $\lambda$ controls the trade-off between reconstruction fidelity and sparsity. The $L_{1}$ penalty encourages most components of $h(x)$ to be close to zero, effectively forcing the autoencoder to represent inputs as a sparse linear combination of a small number of feature directions. This property makes SAEs particularly useful for high-dimensional data, where they can uncover disentangled, interpretable features. In the context of combining SAEs with LLMs and transformers in general, the input $x$ for SAE to reconstruct is typically an embedding extracted either from the transformer’s residual stream or sometimes from MLP or attention layers (Kissane et al., 2024).
@@ 4.1. Training
Our approach for applying SAE to sequential recommendations is illustrated in Figure  1. Firstly, we train a transformer-based model from scratch on the given dataset. The model is trained solely on user-item interactions without incorporating textual or any other side information. This step differs from the NLP domain, where researchers usually analyze a large model pretrained on a huge amount of data. Secondly, we run part of the data through the model and extract activations from some intermediate transformer layer. The activations from each sequence step are considered as a separate data point for training the autoencoder. Next, we train a standard SAE with L1 penalty on these activations as in (Cunningham et al., 2023; Bricken et al., 2023; Templeton et al., 2024).
@@ 4.3. Controlling the Model Behavior
We perform intervention in the forward pass of the model during inference in the following way. We take the activations of the selected transformer layer, encode them with the SAE encoder, set the activation of the selected SAE feature to some desired value, reconstruct the original layer with the SAE decoder, and replace the layer with this reconstruction. If the SAE feature corresponds to some attribute, the model outputs will change accordingly.
@@ 5. Experimental Setting
For our main experiments, we use the well-known and popular Movielens-20m dataset (Harper and Konstan, 2015). It contains sequences of user ratings of movies, as well as their attributes - title, genres, and year. For our analysis we also add information about the movie’s language. This choice of dataset makes it easier to analyze the results since movies and their genres are highly interpretable by humans, and we expect that there should be some preferences regarding genres in user behavior. To ensure the generalizability of the results, we perform additional experiments on the Music4all dataset (Santana et al., 2020), which contains user listening history along with different song attributes, including genres.
@@
As a recommendation model, we use GPTRec (Petrov and Macdonald, 2023), which is based on GPT architecture and is very similar to the well-known SASRec model (Kang and McAuley, 2018). However, the suggested approach can be directly applied to any other choice of transformer-based model. We set the number of transformer layers to 3, the number of attention heads to 2, and the hidden size to 64.
SAE is trained on activations from the output of the first transformer block, as this layer has slightly better interpretability metrics. Nevertheless, selecting other layers leads to similar results. We normalize activations by subtracting the mean and dividing by the standard deviation. Normalization is a common practice (Gao et al., 2024; Rajamanoharan et al., 2024b; Templeton et al., 2024), which makes it easier to transfer hyperparameters and compare different runs on different layers. SAE is trained with Adam optimizer, and the learning rate is 1e-3. We tried different SAE dictionary size values ranging from 64 to 8196 and different weights of L1 penalty ranging from 0.01 to 1. The choice of these hyperparameters is discussed in the following sections.
We use dictionary_learning repository [^1] https://github.com/saprmarks/dictionary_learning for SAE training, NNsight library (Fiotto-Kaufman et al., 2024) for model intervention, and RePlay library (Vasilev et al., 2024) for preprocessing and metrics calculation.
@@ 6. SAE Reconstruction Properties
A sparse autoencoder is fundamentally designed to reconstruct the original activations of the transformer layer. While achieving near-perfect reconstruction is feasible with a sufficiently wide autoencoder and low sparsity, it is not the primary objective – the ultimate goal is to produce an autoencoder with interpretable features. On the other hand, SAE with too low reconstruction quality may not be very useful, as it may poorly reflect the properties of the original model. Another important consideration is the sparsity of the trained SAE. An overly sparse autoencoder with very few active features would struggle to achieve good reconstruction. Conversely, an autoencoder with too many active features risks becoming trivial, potentially losing both interpretability and monosemanticity. Balancing reconstruction quality and sparsity is, therefore, critical for designing an effective SAE.
@@
L1 RMSE Explained L0 Recommendation metrics @10
@@
1 0.922 0.0001 0.5 0.039 0.079 0.103
Table 1 contains root mean squared error (RMSE), a fraction of explained variance, an average number of active features (sparsity), as well as NDCG@10, HitRate@10, and Coverage@10 for different values of L1 regularization and dictionary size of 2048. L1 regularization is the most influential hyperparameter, as small changes in its values have a strong effect on all metrics. For further experiments we set L1 regularization to 0.1, as for this value the reconstruction is good enough, but not perfect. Figure 3 from section 7 confirms that for this value, interpretability is the best.
@@ 7.1. Results
The Movielens-20m dataset contains binary attributes indicating whether the movie belongs to a given genre. We follow the approach described in section 4.2 and compute correlation, ROC AUC, and sensitivity for each SAE feature and genre. For each genre, we find one top feature with the maximum correlation. Figure 2 shows correlations for these top features with all the genres. It is clear that features are rather monosemantic and mostly correlated with one or two similar genres. Animation and Children turned out to be the most interconnected genres since they both share the same top feature. This connection is easily explained because animated movies are often intended for children. The same connection holds for Thriller and Mystery. The Adventure feature has some correlation with Action, the Fantasy feature - with Adventure, the Sci-Fi feature - with Action, and the War feature - with Drama. All these patterns can be explained based on genre meanings.
Table 2. Metrics for SAE features with maximum correlation for each genre and genre popularity (proportion of data points with given genre). The data is sorted by correlation. The last line corresponds to the averages for all genres, which are considered as overall interpretability metrics.
@@
Table 2 contains metric values for these top features. The results vary greatly depending on the genre. For some genres, such as Horror and Sci-Fi, the correlation and ROC AUC are high, and for some others, they are significantly lower. For example, more broad and vague genres like Action, Drama, and Adventure have lower correspondence to features, as they are presented too often and in different contexts.
@@
The last line of Table 2 contains mean metrics across all genres. These values reflect SAE interpretability and can be used as a guideline for selecting SAE hyperparameters. Figure 3 illustrates the dependency of mean correlation on L1 penalty and SAE dictionary size. We make 5 runs for each point and average them, as we found that for a single run, the mean correlation may have some fluctuations. The mean correlation is lower for very low L1 values with perfect reconstruction quality and for high L1 values with bad reconstruction quality (see Table 1). For a dictionary size above 256, the mean correlation is stable.
@@
It is important to check whether SAE features are more interpretable than the neurons of the original transformer layer on which SAE was trained. In order to do this, we perform the same procedure, compute the correlation between neurons of the transformer layer and genres, and find the top neurons with the maximum correlation for each genre. Figure 4(a) contains these correlation values for neurons of the transformer layer along with SAE features. The association with the SAE feature is much higher for all genres except Adventure. Correlation with neurons of the transformer layer is especially low for rare genres, e.g. Documentary, Film-Noir, and Western, while SAE is able to learn good features for these cases. For the transformer layer, the minimum correlation is 0.07, the mean is 0.20, and the maximum is 0.33, while for SAE, the minimum correlation is 0.29, the mean is 0.44, and the maximum is 0.71.
7.1.4. Results on Music4all dataset.
To verify that the results generalize to different datasets, we additionally perform experiments on Music4all dataset and compare SAE features with musical genres. The genre taxonomy is rather messy, as there are 853 unique values. We take 20 most popular genres and find top SAE features and neurons of the original transformer layer, as in the previous section.
Figure 4(b) contains the results, which are quite similar to the Movielens-20m dataset. Association with SAE features is significantly higher for most of the genres, with some genres being more successful than others. For pop music, the correlation with the SAE feature is low, while the correlation with the original neuron is much higher. This is because it is the most popular genre, so it is reflected in the model, while for SAE, it is too broad, so the features capture only its subsets. For the transformer layer, the minimum correlation is 0.09, the mean is 0.16, and the maximum is 0.4, while for SAE, the minimum correlation is 0.14, the mean is 0.32, and the maximum is 0.9.

### v2

@@ Abstract.
Many current state-of-the-art models for sequential recommendations are based on transformer architectures. Interpretation and explanation of such black box models is an important research question, as a better understanding of their internals can help understand, influence, and control their behavior, which is very important in a variety of real-world applications. Recently, sparse autoencoders (SAE) have been shown to be a promising unsupervised approach to extract interpretable features from neural networks.
In this work, we extend SAE to sequential recommender systems and propose a framework for interpreting and controlling model representations. We show that this approach can be successfully applied to the transformer trained on a sequential recommendation task: directions learned in such an unsupervised regime turn out to be more interpretable and monosemantic than the original hidden state dimensions. Further, we demonstrate a straightforward way to effectively and flexibly control the model’s behavior, giving developers and users of recommendation systems the ability to adjust their recommendations to various custom scenarios and contexts.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The International World Wide Web Conference; April 13–17, 2026; Dubai, UAEisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Sequential recommender systems capture the dynamic nature of user preferences by taking into account the temporal order of user actions. This approach is particularly valuable in online environments, such as e-commerce platforms, social networks, and content streaming services, where user interests evolve and change rapidly, and the sequence of interactions strongly influences future behavior. By incorporating this sequential information, models can provide more accurate and relevant recommendations, improving personalization and user experience.
@@
Transformer-based models, such as SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019), and their numerous improvements (Mezentsev et al., 2024; Petrov and Macdonald, 2023b; Tikhonovich et al., 2025; Li et al., 2020; Frolov et al., 2024), effectively model complex sequential patterns. Interpretation and explanation of these models are of high practical importance due to their increasing usage both in research and industrial applications. Explainable recommendations, providing transparency into why certain recommendations are made, can help detect biases, improve user trust and the reliability of recommendations. In addition, interpretability may help to adjust recommendations to specific contexts, individual user preferences, and needs, leading to a more satisfying and personalized experience.
Sparse Autoencoders (SAE) have recently emerged as a powerful unsupervised approach for discovering interpretable directions in the latent representations of neural networks. While this technique has been primarily explored in the domains of language and vision (Cunningham et al., 2023; Makelov et al., 2024; Gao et al., 2024; Gorton, 2024), its application to recommendation models is still at an early stage. Recent work (Wang et al., 2024) has begun to explore the use of sparse autoencoders for interpretation of recommender systems. We further develop SAE-based interpretation methods and conduct a systematic study of the ability to influence model outputs. Recommendation data is structured around user–item interactions rather than tokens, and items are often associated with rich metadata and attributes. These properties require adapting the SAE pipeline to account for domain-specific structure and to define meaningful interpretability metrics grounded in item attributes.
We extend SAE to sequential recommendations and propose a principled methodology for evaluating and interpreting the learned representations. We introduce quantitative metrics for assessing the interpretability of SAE features based on explicit item attributes, as well as procedures for identifying meaningful directions in the model’s latent space. Beyond interpretability, we show that these directions enable precise and flexible control over model behavior by amplifying or suppressing specific item properties - an effect known as steering (Bricken et al., 2023). We perform a systematic analysis demonstrating that steering works reliably in recommendation models and measure its impact on both the semantics and quality of recommendations. Finally, we validate the approach in two distinct domains (music and movies), showing that it generalizes across datasets with different attribute structures and is applicable to a wide range of personalization scenarios.
@@
  • We extend the sparse autoencoders to sequential recommendation models, proposing methods to identify interpretable features and metrics to assess interpretability using item attributes.
  • We demonstrate that SAE learns numerous interpretable, non-trivial, and semantically distinct features that align with meaningful item properties.
  • We show that SAE enables flexible control over model predictions and quantitatively evaluate its effectiveness and influence on recommendations.
@@ 2. Related Work
Various approaches have been proposed to interpret recommendation models. In the study (Tsang et al., 2020), automatic detection of important feature interactions in neural networks is introduced. Other interpretation techniques are based on language modeling (Hada et al., 2021; Li et al., 2021, 2023b), attention mechanism (Xiao et al., 2017; Tal et al., 2021), counterfactual methods (Tan et al., 2021; Kaffes et al., 2021), and other explanation strategies. While these methods help reveal aspects of the model’s internal structure, they are often closely tied to specific architectures, which limits their generalizability. Recent work (Wang et al., 2024) applied SAE for interpreting recommendation models, relying on automated interpretation with a language model based on textual descriptions, an approach inherited from NLP. In contrast, we propose a domain-specific methodology that leverages explicit item attributes to interpret learned representations, introduce new quantitative metrics for evaluating interpretability, and provide a comprehensive analysis of model control.
@@
Some recent works have also applied sparse autoencoders to recommender systems, but with different objectives. In (Kasalickỳ et al., 2025), SAE is used primarily as a compression mechanism, projecting dense embeddings into a high-dimensional, sparsely activated space. Another study (Ahmadov and Mansoury, 2025) employs SAE to identify neurons encoding popularity-related signals and adjust their activations to mitigate popularity bias in a sequential recommendation model.
@@ 3. Sparse Autoencoders
SAE has emerged as a tool in the field of LLM mechanistic interpretability through the employment of techniques from sparse dictionary learning that were previously applied to tasks such as understanding neuronal signals in the visual cortex  (Olshausen and Field, 1997, 2004; Cunningham et al., 2023). Building on experiments with smaller transformer models (Elhage et al., 2022; Cunningham et al., 2023), SAE-based approaches have been further developed and scaled to larger models, including cutting-edge proprietary LLMs. These advancements continue to provide insights into model internals and offer promising strategies for better controlling model generation (Gao et al., 2024).
Sparse autoencoders (SAEs) are a class of autoencoders designed to produce compact and interpretable latent representations by ensuring that only a small subset of hidden units becomes active for any given input. We consider the classic formulation, where an input $x\in\mathbb{R}^{n}$ is mapped to a hidden representation $h(x)\in\mathbb{R}^{m}$ through an encoder and then reconstructed by a decoder to approximate the original input, $\hat{x}$. This can be expressed as $h(x)=ReLU(Wx+b)$ and $\hat{x}=W^{\prime}h(x)+b^{\prime}$, where $W,W^{\prime}$ are the weight matrices, $b,b^{\prime}$ are bias terms. The training objective is to minimize the reconstruction loss, $\|x-\hat{x}\|_{2}^{2}$ while imposing a sparsity constraint on the hidden representation $h(x)$.
A common approach to enforce sparsity on SAE is to add an $L_{1}$ regularization term on $h(x)$, resulting in the following objective function  (Sharkey et al., 2022; Rajamanoharan et al., 2024a): $L(x)=\|x-\hat{x}\|_{2}^{2}+\lambda\|h(x)\|_{1}$, where $\lambda$ controls the trade-off between reconstruction fidelity and sparsity. The $L_{1}$ penalty encourages most components of $h(x)$ to be close to zero, effectively forcing the autoencoder to represent inputs as a sparse linear combination of a small number of feature directions. This property makes SAEs particularly useful for high-dimensional data, where they can uncover disentangled, interpretable features. In the context of combining SAEs with transformers, the input $x$ for SAE to reconstruct is typically an embedding extracted either from the transformer’s residual stream or sometimes from MLP or attention layers (Kissane et al., 2024).
@@ 4.1. Training
Our approach for applying SAE to sequential recommendations is illustrated in Figure  1. Firstly, we train a transformer-based model from scratch on the given dataset. The model is trained solely on user-item interactions without incorporating textual or any other side information. This step differs from the NLP domain, where researchers typically analyze a large model pretrained on a huge amount of data. Secondly, we run part of the data through the model and extract activations from some intermediate transformer layer. The activations from each sequence step are treated as separate data points for training the autoencoder. Next, we train SAE described in Section 3 on these activations.
@@ 4.3. Controlling the Model Behavior
We perform intervention in the forward pass of the model during inference in the following way. We take the activations of the selected transformer layer, encode them with the SAE encoder, set the activation of the selected SAE feature to some desired value, reconstruct the original layer with the SAE decoder, and replace the layer with this reconstruction. This approach is commonly referred to as steering (Bricken et al., 2023). If the SAE feature corresponds to some attribute, the model outputs will change accordingly.
@@ 5. Experimental Setting
Experiments are conducted on the Movielens-20M (Harper and Konstan, 2015) and Music4All (Santana et al., 2020) datasets. The movie and music domains are well-suited for studying model interpretability and control, as user preferences over content categories like genres are often meaningful and diverse. The ability to guide recommendations in such settings can be valuable for improving user experience and personalization. Movielens-20M contains user ratings of movies, as well as their attributes: title, genres, and release year. We additionally include the movie’s language in our analysis. Music4All contains user listening history along with various song attributes, including genres. The genre taxonomy is rather messy, as there are 853 unique values. So, we take 20 most popular genres in the analysis.
@@
We use two transformer-based models: GPTRec (Petrov and Macdonald, 2023a), which follows the transformer decoder architecture and is very similar to the widely-used SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019), which is based on the transformer encoder architecture. This allows us to cover both major classes of transformer-based models used in sequential recommendations. We set the number of transformer layers to 3, the number of attention heads to 2, and the hidden size to 64.
SAE is trained on activations from the output of the first transformer block, as this layer has slightly better interpretability metrics. Nevertheless, selecting other layers leads to similar results. We normalize activations by subtracting the mean and dividing by the standard deviation. Normalization is a common practice (Gao et al., 2024; Rajamanoharan et al., 2024b), which makes it easier to transfer hyperparameters and compare different runs on different layers. SAE is trained with Adam optimizer, and the learning rate is 1e-3. We tried different SAE dictionary size values ranging from 64 to 8196 and different weights of L1 penalty ranging from 0.01 to 1. The choice of these hyperparameters is discussed in the following sections.
Since the results are consistent across models and datasets, we report results for GPTRec on the MovieLens-20M dataset by default, unless stated otherwise. The appendix contains additional results for BERT4Rec and the Music4All dataset. We use dictionary_learning repository [^1] https://github.com/saprmarks/dictionary_learning for SAE training, NNsight library (Fiotto-Kaufman et al., 2024) for model intervention, and RePlay library (Vasilev et al., 2024) for preprocessing and metrics calculation.
@@ 6. SAE Reconstruction Properties
A sparse autoencoder is designed to reconstruct the activations of a transformer layer, but perfect reconstruction is not the main goal – the key objective is to learn interpretable features. If reconstruction quality is too low, SAE may fail to capture meaningful properties of the original model. At the same time, sparsity plays a crucial role: an overly sparse model may lack the capacity for accurate reconstruction, while too many active features can make the representation trivial and reduce interpretability. Therefore, effective SAE design requires balancing reconstruction quality and sparsity.
@@
L1 RMSE Explained L0 Recommend. metrics @10
@@
1.0 0.922 0.0001 0.5 0.039 0.079 0.103
Table 1 contains root mean squared error (RMSE), a fraction of explained variance, an average number of active features (sparsity), as well as NDCG@10, HitRate@10, and Coverage@10 for different values of L1 regularization and dictionary size of 2048. L1 regularization is the most influential hyperparameter, as small changes in its values have a strong effect on all metrics. For further experiments we set L1 regularization to 0.1, as for this value the reconstruction is good enough, but not perfect. Figure 3 from Section 7.1.2 confirms that for this value, interpretability is the best.
@@ 7. Interpretation
We focus our interpretability metrics on genres, as they are clear, well-understood, and widely used in practice. Enabling control over genre-level attributes can be particularly relevant for real-world applications. Additionally, we provide interpretability evaluation using other item attributes in Appendix C.
@@ 7.1. Results
We follow the approach described in Section 4.2 and compute correlation, ROC AUC, and sensitivity between each SAE feature and genre. For each genre, we identify the top feature with the highest correlation. Figure 2 shows correlations of these top features with all genres on the Movielens-20M dataset. The features are mostly monosemantic and correlate with one or two similar genres. Animation and Children are the most interconnected genres since they both share the same top feature, which is expected as animated movies are often intended for children. The same connection holds for Thriller and Mystery. Other patterns (e.g., Adventure feature correlates with Action, Fantasy with Adventure, War with Drama) are also consistent with genre semantics.
Table 2. Metrics for SAE features with maximum correlation for each genre and genre popularity (proportion of data points with given genre). The data is sorted by correlation. The last row shows the averages across all genres, which are considered as overall interpretability metrics.
@@
Table 2 reports interpretability metrics for these top features. The results vary greatly depending on the genre. For some genres, such as Horror and Sci-Fi, the correlation and ROC AUC are high, whereas for others they are significantly lower. For example, broader and more loosely defined genres like Action, Drama, and Adventure have lower correspondence with features, likely because they appear too frequently and across different contexts.
The results for BERT4Rec and the Music4All dataset are similar and provided in the appendix.
@@
The last line of Table 2 contains mean metrics across all genres. These values reflect SAE interpretability and can be used as a guideline for selecting SAE hyperparameters. Figure 3 illustrates the dependency of mean correlation on L1 penalty and SAE dictionary size. We make 5 runs for each point and average them, as we found that the mean correlation can fluctuate across runs. The mean correlation is lower for very low L1 values with perfect reconstruction quality and for high L1 values with bad reconstruction quality (see Table 1). For a dictionary size above 256, the mean correlation is stable.
@@
It is important to check whether SAE features are more interpretable than the neurons of the original transformer layer on which SAE was trained. In order to do this, we perform the same procedure, compute the correlation between neurons of the transformer layer and genres, and find the top neurons with the maximum correlation for each genre.
Figure 4(a) contains these correlation values for top neurons of the transformer layer along with top SAE features for the Movielens-20M dataset. The association with the SAE feature is much higher for all genres except Adventure. Correlation with neurons of the transformer layer is especially low for rare genres, e.g. Documentary, Film-Noir, and Western, while SAE is able to learn good features for these cases.
Table 3. Comparison of SAE features with original transformer layer neurons. Minimum, mean, and maximum correlation values for top features/neurons.
Dataset Model Transformer Layer SAE Features
min mean max min mean max
Movielens-20M GPTRec 0.07 0.20 0.33 0.29 0.44 0.71
Movielens-20M BERT4Rec 0.04 0.14 0.22 0.16 0.29 0.53
Music4all GPTRec 0.09 0.16 0.40 0.14 0.32 0.90
Figure 4(b) contains the correlation values for the Music4all dataset. Association with SAE features is significantly higher for most of the genres, with some genres being more successful than others. For pop music, the correlation with the SAE feature is low, while the correlation with the original neuron is much higher. This is because it is the most popular genre, which is reflected in the model, whereas for SAE it is too broad, so the features capture only its subsets.
The corresponding figure for BERT4Rec is provided in Appendix A. Table 3 summarizes the results, showing the minimum, mean, and maximum correlation for both the neurons of the transformer layer and the SAE features. Although the absolute values for SAE features for BERT4Rec are lower than those for GPTRec, this is due to weaker genre alignment in the original BERT4Rec layer. Nevertheless, SAE still clearly improves interpretability.